OpenStudy (anonymous):

What is the value of integral 0 to 4 3x^2 sqrt (x^3 + 1) dx?

OpenStudy (fibonaccichick666):

well, what do you think you can do?

OpenStudy (anonymous):

That's sort of why I'm on here...to get a little help.

OpenStudy (anonymous):

Hint\[ \int_0^4 3x^2\sqrt{x^3+1}~dx = \int_0^4\sqrt{x^2+1}~d(x^3+1) \]

OpenStudy (solomonzelman):

well u=x^3+1

OpenStudy (solomonzelman):

u = x^3+1, so your du is going to be equal to dx times the derivative of u.

OpenStudy (fibonaccichick666):

so, what happens if you do a substitution?

OpenStudy (fibonaccichick666):

like the one solomon suggested

OpenStudy (anonymous):

Ok so du = dx(3x^2), right?

OpenStudy (solomonzelman):

yes du=3x^2 dx

OpenStudy (fibonaccichick666):

yep, so now, can you integrate your remaining function?

OpenStudy (solomonzelman):